Please help
A. Find the area bounded between the curves : 1.) F(x) = x and G(x) = 2x - x and x = -2 and y= 0.
Answer #1

9=23 psy=2x-x² Area = q x3-12x-x²) du = kosa ( 4x4 - 4-8)